OpenStudy (anonymous):

Evaluate (5/7 - 4/9) * 2/3 + (-1 1/4)

OpenStudy (anonymous):

It can be represented as A*B + C A = \[\frac{5}{7} - \frac{4}{9}\] B = \[\frac{2}{3} \\ C = \frac{-3}{4}\] A = \[\frac{(5*9 - 4*7)}{7*9} = \frac{45 - 28}{63} = \frac{17}{63}\] \[A*B = \frac{17*2}{63*3} = \frac{34}{189}\] \[A*B + C = \frac{34*4 + 189*(-3)}{189*4} = \frac{-431}{756}\]

OpenStudy (anonymous):

What do u think the answer is

OpenStudy (mathwizzard3):

It looks like someone already explained